OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of ?

OpenStudy (zzr0ck3r):

has a slope of what?

OpenStudy (anonymous):

2/3

OpenStudy (zzr0ck3r):

ok the equation of a line is y=mx+b where m is the slope and b is the y intercept so we know y = (2/3)x+b we need to know what b is so we plug in the point they gave us (6,3) let x = 6, and y = 3 y=(2/3)x+b becomes 3=(2/3)(6)+b 3=4+b b = -1 so \[y=\frac{2}{3}x-1\]

OpenStudy (anonymous):

yea but this has to be in standard form..

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (luigi0210):

@zzr0ck3r Was close. Standard form is Ax+By=C, where A is not a negative or fraction answer

OpenStudy (anonymous):

yea.. can you help me figure it out??

OpenStudy (luigi0210):

Pretty simple. Just move the x over and get it to be 1x. I'm not here to give answers or do your work for you though, which @zzr0ck3r did and shouldn't really be doing.

OpenStudy (zzr0ck3r):

I showed my steps, I had to go...you all were not helping. I got a bunch of messages from the user.....again you all were not helping...@Luigi0210 I did not just give an answer, i got him part of the way there. I hope he would be able to do it again if asked.

OpenStudy (anonymous):

yea..but IDK how to set it up

OpenStudy (anonymous):

do you?

OpenStudy (anonymous):

@Luigi0210 CAn you help me?